OpenStudy (anonymous):

The sum of three numbers is 137. The second number is four more than two times the first number. The Third number is five less than three times the first number. find the three numbers.

OpenStudy (anonymous):

x+y+z=137 y=2x+4 z=3x-5 solve these and you are done, do you need help?

OpenStudy (anonymous):

let first no = x then second no = 2x + 4 third number = 3x -5 so x + (2x+4) + (3x-5) = 137 => 6x - 1 = 137 => 6x = 137+1 => x = 138/6 => x = 23 so first no is 23 second no is 2(23) + 4 = 50 third no is 3(23) - 5 = 64